The sum of the digits of a two-digit number is one third of the number. The digits unit is five more than the tens digit. What is the number?
Answer by checkley79(3341)

x+y=(10x+y)/3
y=x+5
x+(x+5)=(10x+x+5)/3
2x+5=(11x+5)/3
3(2x+5)=11x+5
6x+15=11x+5
6x-11x=5-15
-5x=-10
x=-10/-5
x=2 ans. for the tens digit.
y=2+5=7 ans. for the units digit.
Proof:
2+7=(10*2+7)/3
9=(20+7)/3
9=27/3
9=9